OpenStudy (zuha):

2. Find the slope of a line that passes through (–2, –3) and (1, 1). 1/3 1 2 4/3

OpenStudy (bradely):

m = (y2 - y1)/(x2 - x1) Here, x1 = -2,x2 = 1, y1 = -3, and y2 = 1 m = [1-(-3)]/[1- (-2)] = 4/3 Source: http://www.mathskey.com/question2answer/

OpenStudy (zuha):

Thanks